Given that: C=$40m, I=$60m, G=$20m, T=$20m.

Find :

1. Equilibrium Level of income.

2. Consumption expenditure.

3. Level of savings.

4. New income Level if investment falls by $10m.


1
Expert's answer
2021-03-26T08:13:26-0400

1. Equilibrium Level of income is:

Y = C + G + I = 40 + 20 + 60 = 120.

2. Consumption expenditure is:

C = 40.

3. Level of savings is:

S = Y - C - G = 60.

4. If investment falls by $10m, then:

Y = 40 + 20 + 50 = 110.